# Gauge-Mediated Contagion: A Quantum Electrodynamics-Inspired Framework for Non-Local Epidemic Dynamics and Superdiffusion

### v1

@@ X Conclusion and Future Perspectives
In this work, we have proposed a QED-inspired $SIR$ model. In this model, all the dynamics of the disease transmission is implicitly defined. By replacing the ”direct contact” paradigm of classical SIR models with a dynamic interaction mediated by a mediator field $\varphi$, we have demonstrated that QED mathematical techniques can be extended to this epidemiological model giving us mechanisms to describe various observables than the standard $SIR$ model fails to explain as non-local interactions, non-linear effects,the onset of a pandemic and variation of $R_{eff}$. Crucially, we show that the non-local interactions observed in real-world data are not ad-hoc additions without justifications more than to fit the data but emerge naturally from the theory by integrating out the pathogen field. We also have shown how naturally spatial correlations are emerging.
@@
We have identified the epidemic threshold not only as a statistical parameter ($R_{0}$), but as a symmetry-breaking phase transition where the vacuum becomes unstable. This formalism offers a theoretical basis for “Pandemic Opalescence”, explaining the divergence of correlation lengths and the critical fluctuations observed prior to major outbreaks.

### v2

@@ XI Conclusion
In this work, we have proposed a QED-inspired SIRSIR model. In this model, all the dynamics of the disease transmission is implicitly defined. By replacing the ”direct contact” paradigm of classical SIR models with a dynamic interaction mediated by a mediator field $\varphi$, we have demonstrated that QED mathematical techniques can be extended to this epidemiological model giving us mechanisms to describe various observables than the standard SIRSIR model fails to explain as non-local interactions, non-linear effects,the onset of a pandemic and variation of $R_{eff}$. Crucially, we show that the non-local interactions observed in real-world data are not ad-hoc additions without justifications more than to fit the data but emerge naturally from the theory by integrating out the pathogen field. We also have shown how naturally spatial correlations are emerging.
@@
We have identified the epidemic threshold not only as a statistical parameter ($R_{0}$), but as a symmetry-breaking phase transition where the vacuum becomes unstable. This formalism offers a theoretical basis for “Pandemic Opalescence”, explaining the divergence of correlation lengths and the critical fluctuations observed prior to major outbreaks. The “epidemiological vacuum”—the structural readiness of a network to sustain widespread contagion—is quantified by the effective mass parameter, $m_{R}(t)$. We demonstrated that the approach to a macroscopic outbreak is reliably preceded by the collapse of this effective mass ($m_{R}(t)\to 0$), which corresponds to a divergence in the structural correlation length of the network.
These results have been validated using spatial data from the COVID-19 pandemic in Germany. The gauge field serves as a sensitive leading indicator. The collapse of $m_{R}(t)$ anticipated localized surges in empirical case counts, significantly outperforming classical reactive metrics such as the effective reproduction number ($R_{eff}$).
By linking the formalism of quantum field theory with network epidemiology, this work provides a paradigm shift from reactive observation to structural prediction. The ability to monitor the “tension” of the epidemiological field in real-time offers public health authorities a new tool for early warning and targeted intervention.
@@
Acknowledgements.
We acknowledge financial support from SECIHTI and SNII (México) and Guanajuato University.